OpenStudy (anonymous):

find the slope of the line through each pair of points (3/2,-1/2) and (-2/3,1/3)

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

1/3 - (-1/2)/ -2/3 - 3/2

OpenStudy (anonymous):

get it? just find the answer

hero (hero):

Insert each point into y = mx + b as follows: -1/2 = 3/2(m) + b 1/3 = -2/3(m) + b Isolate b in each: -1/2 - 3/2m = b 1/3 + 2/3 m = b set b = b -1/2 - 3/2m = 1/3 + 2/3m Place like terms on both sides: -1/2 - 1/3 = 2/3m + 3/2 m -3/6 - 2/6 = 4/6m + 9/6m -5/6= 13/6m Multiply both sides by 6/13 -5/13 = m

OpenStudy (anonymous):

Remember slope is equal to rise/run so you take the rise or the difference of the y coordinates and divide them by the difference of the x coordinates so you get: \[m=\left( y _{2} - y _{1}\right)\div \left( x _{2}-x _{1}\right)\] \[\left( 1/3-(-1/2) \right)\div \left( (-2/3)-3/2 \right)\] m=-5/13